### v1

@@ Abstract
To holistically evaluate the performance of a quantum metrology scheme, several other important figures of merit must also be considered in conjunction with the maximum FI. Specifically, these include the parameter range in which we achieve quantum-enhanced precision, the corresponding average sensitivity, and the capability to shift the maximum sensitivity point to origin $\theta=0$ which is crucial for sensing applications. We define the bandwidth as the range of $\theta$ in which the FI of SCS is above the maximum FI of the CS, $\tilde{\text{FI}}$ as the average FI within the bandwidth, and $\theta$ offset where the FI is maximum. The bandwidth for SCS up to $\overline{n}=1.76$ is nearly constant at approximately 0.5 rad, which sets the limit on the allowed phase fluctuations in the system. The $\tilde{\text{FI}}$ against photon number $\overline{n}$ in Fig. 3(a) shows a quadratic trend, suggesting that working within the bandwidth still affords favourable precision scaling. The offset in $\theta$ is close to the position of the maximum slope $|\partial p/\partial\theta|$ in the measured probability [33] and decreases for SCS with larger $\alpha$, as shown in Fig. 3(a). With our versatile protocol, this offset is a fully programmable parameter and can be adjusted by simply preparing input states with an initial phase offset (Fig. 3(b)) to ensure that we can operate around the maximum sensitivity point.
@@
In addition to the FI, which only sets the lower bound to the precision according to the Cramér-Rao bound, we now present a benchmark for direct calculation of the phase precision, $\Delta\theta$. It is computed as $\Delta\theta=\Delta p/|\partial p/\partial\theta|$, where $\partial p/\partial\theta$ is obtained from the polynomial fit as in Fig. 2(c) inset and $\Delta p$ is the standard deviation of the single shot measurement with $1000$ repetitions. The standard deviation for SCS (blue circles) and CS (purple diamonds) are plotted in Fig. 3(c), which respectively beat the precision limit $\Delta\theta=1/\overline{n}$ and the SQL ($\Delta\theta=1/\sqrt{\overline{n}}$) commonly used in optical [14, 15] or atomic [45] interferometers. We also note that the experimentally obtained precision for SCS, $\Delta\theta_{\mathrm{SCS}}$, surpasses the ideal limit used for the superposition of Fock states ($\Delta\theta=1/2\overline{n}$[27, 28]. This is because the QFI of SCS and CS in our optimised strategy gives $\Delta\theta=1/\sqrt{4\overline{n}+4\overline{n}^{2}}$ and $\Delta\theta=1/\sqrt{4\overline{n}}$, respectively [33] to afford superior performance by design. Furthermore, the linear fit of the SCS shows that $\Delta\theta_{\text{SCS}}$ approaches the Heisenberg limit scaling. To benchmark the quantum enhancement of our implementation, we compute the metrological gain of $20\log_{10}(\Delta\theta_{\text{CS}}/\Delta\theta_{\text{SCS}})=7.5(6)$ dB with a low photon number of 1.76. Here, the uncertainty is one standard deviation obtained from bootstrapping and stabilises at $\sim 600$ measurement repetitions [33].
@@
Moreover, the gain or bandwidth can be enhanced on demand by optimizing the initial states with respect to system constraints. In typical hardware implementations, a device may allow access to a limited maximum photon number $\overline{n}$ or SCS amplitude $\alpha$, which may be due to photon loss, self-Kerr effect, etc. Depending on the specific hardware limitations and the figures of merit, we can optimise the metrological performance by simply varying the superposition weights of the SCS. Physically, this is achieved by adjusting the qubit and cavity drives for $\hat{O}_{1,2}$, without any changes to the device or system dynamics. For the case where the constraint is the photon number, e.g., $\overline{n}=1$, we show that higher maximum FI can be achieved by putting more weight on the $|0\rangle$ component of the SCS, at the expense of sensitivity range, as shown in Fig. 4(a). On the other hand, if the constraint is given by the maximum amplitude, e.g., $\alpha=1.6$, we show states with a higher weight on $|\alpha\rangle$ give a higher maximum FI, without compromising the sensitivity range, see Fig. 4(b). Essentially, in both examples, states with higher photon number variance result in higher maximum FI. We can also make use of the illustration given in Fig. 2(b). If $\alpha$ is fixed, the Wigner function for states with more weight $w$ on the $|\alpha\rangle$ component will have a stronger blob at $\alpha$. This way the change of probability $p_{g0}$ is from 1 to a number which gets closer to zero with increasing $w$, resulting in higher sensitivity. However, if $\overline{n}$ is fixed, increasing the weight on the $|\alpha\rangle$ component comes at the cost of using a smaller $\alpha$, which is a more dominant factor. As a result, the sensitivity is reduced.
To illustrate the versatility of our protocol, we demonstrate quantum-enhanced precision for amplitude estimation by simply changing the process unitary in Fig. 2(a) to a displacement in the imaginary direction $\hat{U}(\gamma)=\exp(i\gamma(\hat{a}^{\dagger}+\hat{a}))$, where $\gamma$ is the displacement amplitude. The process generator is now $\hat{x}=(\hat{a}^{\dagger}+\hat{a})$, making states with high variance $(\Delta\hat{x})^{2}$ useful for amplitude estimation. Here, SCS offer high variance $\propto\overline{n}$ compared to CS, which have a constant variance [33]. The illustration is similar to that of phase estimation in Fig. 2(b), except the SCS are now displaced in the imaginary direction instead of rotated. SCS with higher $\alpha$ have more and sharper fringes in between the blobs, such that the overlap with the initial state reduces more sharply as $\gamma$ increases.
Experimentally, we show that SCS displaced vertically in the imaginary direction (SCS-v) offer the highest FI, compared to both CS and SCS displaced horizontally in the real direction (SCS-h), as shown in Fig. 4(c). This can be explained by noting that SCS-h have a different process unitary $\hat{U}(\gamma)=\exp(i\gamma i(\hat{a}-\hat{a}^{\dagger}))$, where the process generator is now $\hat{y}=i(\hat{a}-\hat{a}^{\dagger})$, such that the variance of SCS is low, even lower than that of the CS.
Furthermore, by operating at the maximum FI, we show that the SCS-v (CS) exhibits a linear (constant) trend against $\overline{n}$, as shown in Fig. 4(d). This suggests a $\sqrt{\overline{n}}$ quantum-enhanced precision of the SCS-v over the CS. We refer to Ref. [33] for a comprehensive analysis of the FI, which importantly includes the shift of the max FI to the initial position $\gamma=0$. The $\sqrt{\overline{n}}$ enhancement also manifests in the direct calculation of the amplitude precision, which reveals that $\Delta\gamma\propto 1/\sqrt{\overline{n}}$ for SCS-v and $\Delta\gamma\approx$ constant for CS [33]. Finally, directly comparing SCS-v to CS reveals a high metrological gain of $9.3(5)$ dB.
@@
In summary, we have introduced a versatile and on-demand protocol for parameter estimation using a single bosonic mode. With this scheme, we demonstrated both phase and amplitude estimation using the superposition of coherent states and achieved quantum-enhanced precision $\sim\sqrt{\overline{n}}$ in both cases, compared to classically behaving coherent states. Using states with photon numbers of up to only 1.76, the high variance of SCS results in state-of-the-art metrological gain, $7.5(6)$ dB for phase and $9.3(5)$ dB for amplitude estimation. We also demonstrated the on-demand adaption of the scheme to optimise for certain metrological performance metrics with respect to hardware constraints.
@@
The simplicity and effectiveness of our protocol stem from the use of a single bosonic mode with two general state-transfer operations that can be realised by any standard gates. As such, our implementation is accessible to systems with different device parameters. Furthermore, the strategy is universal and can be readily performed with many other bosonic systems beyond cQED using their preferred native operations. Our results demonstrate a simple yet powerful strategy for real metrological applications. For instance, quantities such as stray magnetic, electric field, etc. that may affect the qubit or cavity frequency will result in the change of dispersive coupling strength $\chi$ [46], which in turn leads to phase variations on the bosonic state. Hence, the phase precision $\Delta\theta$ obtained in our implementation is directly related to the sensing precision of the relevant physical parameters under investigation. Therefore, our work offers a promising first step towards practical and optimal quantum metrology using a single bosonic system.
@@
Acknowledgement. This project is supported by the National Research Foundation, Singapore, grant number NRFF12-2020-0063 and NRF2020-NRF-ISF004-3540. RF and KP acknowledge support of 21-13265X of the Czech Science Foundation and EU H2020-WIDESPREAD-2020-5 project NONGAUSS (951737) under the CSA - Coordination and Support Action.
Supplementary Information
I Optimality of the protocol
@@
where the absolute term indepedent of $\beta$ represents QFI.
II device and system parameters
Our device consists of a standard three-dimensional (3D) cavity, an ancillary transmon and a planar readout resonator, see Fig. 5. The cavity is made of high-purity (4N) aluminum and is a $\lambda/4$ cylinder-shaped transmission line surrounded by a wall. The readout resonator is deposited together with the transmon on the sapphire using double-angle evaporation technique. The sapphire is inserted in the tunnel, with the pad slightly extending to the cavity. The cavity serves as a high-Q quantum memory to store a bosonic mode and is designed to dispersively couple with the ancillary transmon qubit, which provides the necessary nonlinearity. The low-Q readout resonator is also dispersively coupled with the transom qubit, enabling fast single-shot measurement.
@@
Cavity frequency (GHz)
$\omega_{c}/2\pi$
4.587
Qubit frequency (GHz)
$\omega_{q}/2\pi$
5.277
Resonator frequency (GHz)
$\omega_{r}/2\pi$
7.617
Cavity single-photon lifetime (ms)
$T_{c,1}$
1
Transmon energy lifetime $(\mu s)$
$T_{q,1}$
96
Transmon coherence lifetime $(\mu s)$
$T_{q,2}$
$15.47$
Cavity anharmonicity (kHz)
$K/2\pi$
$6$
Transmon anharmonicity (kHz)
$\alpha/2\pi$
$175.3$
1st cavity-transmon dispersive shift (MHz)
$\chi/2\pi$
1.423
2nd cavity-transmon dispersive shift (MHz)
$\chi^{\prime}/2\pi$
15.76
Transmon-resonator diespersive shift (MHz)
$\chi_{qt}/2\pi$
0.64
@@
Note that for sufficiently high transmon anharmonicity, it can be effectively treated as a qubit (two-level system with $|g\rangle$ ($|e\rangle$) as the ground (excited) level). The first two terms in the bare Hamiltonian then becomes $\Delta\hat{a}^{\dagger}\hat{a}-\chi\hat{a}^{\dagger}\hat{a}|e\rangle\langle e|$, which reduces to $(\chi/2)\hat{\sigma}_{z}\hat{a}^{\dagger}\hat{a}$ for $\Delta=\chi/2$ that is chosen.
III Preliminary investigations using different states
@@
In the context of phase estimation, we investigate several states such as CS $|\alpha\rangle$, superposition of Fock states $(|0\rangle+|N\rangle)/\sqrt{2}$, squeezed vacuum $e^{\frac{-r}{2}(\hat{a}\hat{a}-\hat{a}^{\dagger}\hat{a}^{\dagger})}|0\rangle$, and SCS $\mathcal{N}(|0\rangle+|\alpha\rangle)$. Here, the normalisation factor $\mathcal{N}_{\alpha}=1/\sqrt{2(1+e^{-|\alpha|^{2}/2})}$. The average photon number $\overline{n}=\langle\hat{a}^{\dagger}\hat{a}\rangle$ of these states are respectively $|\alpha|^{2}$, $N/2$, $\sinh^{2}{r}$, and $\mathcal{N}_{\alpha}^{2}|\alpha|^{2}$. The QFI $4(\Delta\hat{n})^{2}$ of these states in terms of $n$ are respectively $4\overline{n}$, $4\overline{n}^{2}$, $8\overline{n}^{2}+8\overline{n}$, and $4((1-\mathcal{N}_{\alpha}^{2})/\mathcal{N}_{\alpha}^{2})\overline{n}^{2}+4\overline{n}$. The QFI of SCS reduces to $4\overline{n}^{2}+4\overline{n}$ for large $\overline{n}$ as $\mathcal{N}_{\alpha}\approx 1/\sqrt{2}$.
Figure 6: Phase and amplitude estimation utilizing various states. The FI of phase estimation (a) and amplitude estimation (b) utilizing CS, superposition of Fock states, squeezed vacuum (Sqvac) and SCS with fixed average photon numbe $n=0.5$.
@@
Experimentally, we realise these states and compute the corresponding FI with the same method as described in the main text, see Fig. 6(a). The comparison is made for states with the same photon number $\overline{n}=0.5$. It is apparent that the SCS offers the highest FI. Although squeezed vacuum is supposed to give a higher theoretical QFI, the state is harder to generate as it requires high squeezing. Highly squeezed states exhibit increased spreading in phase space, making them susceptible to the Kerr effect, which, in turn, hinders quantum-enhanced sensitivity. A viable strategy to mitigate the impact of the Kerr effect involves minimizing the coupling between the cavity and transmon. By reducing this coupling, the undesirable consequences of the Kerr effect can be alleviated, contributing to enhanced quantum sensitivity. Therefore, we discern that with our device, the optimal states are SCS.
Similarly, in the context of amplitude estimation the QFI $4(\Delta\hat{x})^{2}$ for CS, squeezed vacuum, and SCS are respectively given by $4$, $48\sqrt{\overline{n}(\overline{n}+1)}+8\overline{n}+$ 4, and $48\overline{n}+$ 4. The QFI for superposition of Fock states is $4$, $12+4\sqrt{2}$, and $48\overline{n}+$ 4 for $\overline{n}=0.5$, $\overline{n}=1$, and $\overline{n}>1$ respectively.
We experimentally perform the protocol and compute the FI, see Fig. 6(b). Here, unlike the phase estimation, the highest FI is achieved by squeezed vacuum followed by SCS. However, squeezed vacuum with higher $n$ is less robust to Kerr and harder to realise. Therefore, we proceed with SCS for more detailed analysis.
The phase and amplitude estimation underscore the importance of selecting appropriate quantum states to achieve the best precision. Accompanied with practical considerations, this presents an optimal quantum metrology.
IV Imperfection of the system
@@
Figure 8: Performance of quantum-enhanced amplitude precision. (a) The average FI of SCS for which it is above the max FI of CS and $\gamma$ offset at which FI is maximum vs $n$. The dotted line is lineaer fit of $3.2(2)\overline{n}^{1.0(3)}+2.2(1)$. (b) The same protocol as in Fig. 4(c) in main text with initial states prepared with an offset. (c) The amplitude precision vs $n$ for both SCS (blue circles) and CS (purple diamonds). The dotted lines are linear fit of $\log(\Delta\gamma)=a\log(\Delta n)+b$ with $a$ and $b$ as free parameters. The fitted parameters are $a,b=-0.460(3),-1.311(2)$ for SCS and $a,b=-0.0105(2),-0.480(1)$ for CS.
In the experiment, we use gradient ascent pulse engineering (GRAPE) pulse to realize the two state transfer of the protocol. The GRAPE pulses ($\zeta(t)$ and $\epsilon(t)$) are optimized in an ideal system without imperfections. In practice, each element of the system is susceptible to environmental noise, making it crucial to account for the coherence and energy decay rate of both the cavity and transmon qubit. Table 1 presents the corresponding parameters. To account for decoherence, the simulations in the main text (all the sequence) are performed with the following master equation:
Figure 9: Demonstration of amplitude estimation by state optimization. The FI of SCS of different weights with (a) fixed photon number $n=1$ and (b) fixed $\alpha=1.6$. The corresponding shaded area indicates standard deviation from bootstrap.

### v2

@@ IV METROLOGICAL FIGURES OF MERIT
For a holistic evaluation of the performance of a quantum metrology scheme, several other important figures of merit must also be considered in conjunction with the maximum FI. Specifically, these include the parameter range in which we achieve quantum-enhanced precision, the corresponding average sensitivity, and the capability to shift the maximum sensitivity point to origin $\theta=0$ which is crucial for sensing applications. We define the dynamical range as the range of $\theta$ in which the FI of SCS is above the maximum FI of the CS, $\tilde{\text{FI}}$ as the average FI within the dynamical range, and $\theta$ offset where the FI is maximum. The dynamical range for SCS up to $\overline{n}=1.76$ is nearly constant at approximately 0.5 rad, which sets the limit on the allowed phase fluctuations in the system. The $\tilde{\text{FI}}$ against photon number $\overline{n}$ in Fig. 3(a) shows a quadratic trend, suggesting that working within the dynamical range still affords favorable precision scaling. The offset in $\theta$ is close to the position of the maximum slope $|\partial p/\partial\theta|$ in the measured probability and decreases for SCS with larger $\alpha$, as shown in Fig. 3(a). With our versatile protocol, this offset is a fully programmable parameter and can be adjusted by simply preparing input states with an initial phase offset (Fig. 3(b)) to ensure that we can operate around the maximum sensitivity point.
In addition to the FI, which only sets the lower bound to the precision according to the Cramér-Rao bound, we now present a benchmark for direct calculation of the phase precision, $\Delta\theta$. It is computed as $\Delta\theta=\Delta p/|\partial p/\partial\theta|$, where $\partial p/\partial\theta$ is obtained from the polynomial fit as in Fig. 2(c) inset and $\Delta p$ is the standard deviation of the single shot measurement with $1000$ repetitions. The standard deviation for SCS (blue circles) is plotted in Fig. 3(c). The ideal precision, $\Delta\theta=1/\sqrt{4\overline{n}+4\overline{n}^{2}}$ (see Appendix C), surpasses the precision scaling $\Delta\theta=1/\overline{n}$ commonly used in optical [14, 15] or atomic [34] interferometers, as well as that achieved using superposition of Fock states ($\Delta\theta=1/2\overline{n}$[27, 28]. A linear fit of the results obtained using SCS also shows that $\Delta\theta_{\text{SCS}}\propto\bar{n}^{-0.915}$ in the range $\bar{n}=[0.07,1.76]$, which approaches the Heisenberg limit scaling $\Delta\theta\propto\bar{n}^{-1}$. To benchmark the quantum enhancement of our implementation, we compute the metrological gain of $20\log_{10}(\Delta\theta_{\text{CS}}/\Delta\theta_{\text{SCS}})=7.5(6)$ dB with a low photon number of 1.76. Here, the uncertainty is one standard deviation obtained from bootstrapping and stabilizes at $\sim 600$ measurement repetitions, see Appendix E. With only minor adjustments to the protocol, we also showcase amplitude estimation with significant quantum enhancement, see Appendix E.
@@
V ON DEMAND OPTIMISATION SUBJECTED TO SYSTEM CONSTRAINTS
@@
Aside from the significant quantum enhancement achieved, our implementation also has the unique versatility of on-demand optimization for the specific hardware constraints present. Concretely, we show that the gain or dynamical range can be further enhanced without any hardware adjustment by optimizing the weights of the SCS. For instance, typical hardware implementations may often allow access to a limited maximum photon number $\overline{n}$ before the errors due to photon loss or the self-Kerr effect become too prominent. Physically, this is achieved by adjusting the qubit and cavity drives for $\hat{O}_{1,2}$, without any changes to the device or system dynamics. For the case where the constraint is the photon number, e.g., $\overline{n}=1$, we show that higher maximum FI can be achieved by putting more weight on the $|0\rangle$ component of the SCS, at the expense of sensitivity range, as shown in Fig. 4(b) and (c). Intuitively, if $\overline{n}$ is fixed, increasing the weight on the $|\alpha\rangle$ component comes at the cost of using a smaller $\alpha$, which leads to a reduction in the sensitivity. This trade-off between sensitivity and dynamical range as a function of the weights of the SCS is shown in Fig. 4(c). This ability to perform such on-demand optimization is an intrinsic and powerful feature of our implementation, highlighting its practical feasibility for performing quantum metrology in the presence of realistic physical limitations in the hardware.
VI CONCLUSIONS
@@
In summary, we have introduced a practical and versatile protocol for parameter estimation using a single bosonic mode. With this scheme, we demonstrated phase as well as amplitude estimation using the superposition of coherent states. Using states with average photon numbers of up to only 1.76, the high variance and robustness of SCS result in state-of-the-art metrological gain, $7.5(6)$ dB for phase and $9.3(5)$ dB for amplitude estimation, see Appendix E. We also demonstrated the on-demand adaption of the scheme, utilizing SCS with different weights, to optimize for certain metrological performance metrics with respect to hardware constraints.
@@
The efficiency and versatility of our protocol stem from the use of a single bosonic mode with two state-transfer operations that can be realized by any standard gates. As such, our implementation is accessible to systems with different device parameters. Furthermore, the strategy is adaptable and can be readily performed with many other bosonic systems beyond cQED using their preferred native operations. Our results demonstrate a simple yet powerful strategy for real metrological applications. For instance, the phase estimation strategy can be adapted to directly estimate the coupling strength $\chi$ via Bayesian inference (see Appendix F). Future experiments may also extend this direction to estimate other quantities such as stray magnetic, electric field, etc. that may affect the qubit or cavity frequency, which in turn result in the change of dispersive coupling strength $\chi$ [45]. Therefore, our work offers a promising first step towards practical and optimal quantum metrology using a single bosonic system.
ACKNOWLEDGMENT
@@
This project is supported by the National Research Foundation, Singapore, grant number NRFF12-2020-0063 and NRF2020-NRF-ISF004-3540. R.F. acknowledges the support of the Czech Science Foundation via Grant number 21-13265X. K.P. acknowledges the project CZ.02.01.01/00/22__\__008/0004649 (QUEENTEC) of EU and MEYS Czech Republic.
Appendix A Optimality of the protocol
@@
where the absolute term independent of $\beta$ represents QFI.
@@
Appendix B device and system parameters
Our device consists of a standard three-dimensional (3D) cavity, an ancillary transmon and a planar readout resonator, see Fig. 5. The cavity is made of high-purity (4N) aluminum and is a $\lambda/4$ cylinder-shaped transmission line surrounded by a wall. The readout resonator is deposited together with the transmon on the sapphire using a double-angle evaporation technique. The sapphire is inserted in the tunnel, with the pad slightly extending to the cavity. The cavity serves as a high-Q quantum memory to store a bosonic mode and is designed to dispersively couple with the ancillary transmon qubit, which provides the necessary nonlinearity. The low-Q readout resonator is also dispersively coupled with the transom qubit, enabling fast single-shot measurement.
@@
Cavity frequency (GHz) $\omega_{c}/2\pi$ 4.587
Qubit frequency (GHz) $\omega_{q}/2\pi$ 5.277
Resonator frequency (GHz) $\omega_{r}/2\pi$ 7.617
Cavity single-photon lifetime (ms) $T_{c,1}$ 1
Transmon energy lifetime $(\mu s)$ $T_{q,1}$ 96
Transmon coherence lifetime $(\mu s)$ $T_{q,2}$ $15.47$
Cavity anharmonicity (kHz) $K/2\pi$ $6$
Transmon anharmonicity (kHz) $\alpha/2\pi$ $175.3$
1st cavity-transmon dispersive shift (MHz) $\chi/2\pi$ 1.423
2nd cavity-transmon dispersive shift (MHz) $\chi^{\prime}/2\pi$ 15.76
Transmon-resonator dispersive shift (MHz) $\chi_{qt}/2\pi$ 0.64
@@
Figure 6: Phase estimation utilizing various states. The FI of phase estimation utilizing CS, superposition of Fock states, squeezed vacuum (Sqvac) and SCS with fixed average photon number $\bar{n}=0.5$.
Note that for sufficiently high anharmonicity, the transmon can be effectively treated as a qubit (two-level system with $|g\rangle$ ($|e\rangle$) as the ground (excited) level). The first two terms in the bare Hamiltonian then becomes $\Delta\hat{a}^{\dagger}\hat{a}-\chi\hat{a}^{\dagger}\hat{a}|e\rangle\langle e|$, which reduces to $(\chi/2)\hat{\sigma}_{z}\hat{a}^{\dagger}\hat{a}$ for $\Delta=\chi/2$ that is chosen.
Appendix C Preliminary investigations using different states
@@
In the context of phase estimation, we investigate several states such as CS $|\alpha\rangle$, superposition of Fock states $(|0\rangle+|N\rangle)/\sqrt{2}$, squeezed vacuum $e^{\frac{-r}{2}(\hat{a}\hat{a}-\hat{a}^{\dagger}\hat{a}^{\dagger})}|0\rangle$, and SCS $\mathcal{N}(|0\rangle+|\alpha\rangle)$. Here, the normalisation factor $\mathcal{N}_{\alpha}=1/\sqrt{2(1+e^{-|\alpha|^{2}/2})}$. The average photon number $\overline{n}=\langle\hat{a}^{\dagger}\hat{a}\rangle$ of these states are respectively $|\alpha|^{2}$, $N/2$, $\sinh^{2}{r}$, and $\mathcal{N}_{\alpha}^{2}|\alpha|^{2}$. The QFI $4(\Delta\hat{n})^{2}$ of these states in terms of $\bar{n}$ are respectively $4\overline{n}$, $4\overline{n}^{2}$, $8\overline{n}^{2}+8\overline{n}$, and $4((1-\mathcal{N}_{\alpha}^{2})/\mathcal{N}_{\alpha}^{2})\overline{n}^{2}+4\overline{n}$. The QFI of SCS reduces to $4\overline{n}^{2}+4\overline{n}$ for large $\overline{n}$ as $\mathcal{N}_{\alpha}\approx 1/\sqrt{2}$.
Experimentally, we realize these states and compute the corresponding FI with the same method as described in the main text, using Eq. (4), see Fig. 6. The comparison is made for states with the same photon number $\overline{n}=0.5$. We also computed the Fisher information using four probabilities by considering the qubit measurement outcome in both ground and excited states: $p_{g0}(\theta),1-p_{g0}(\theta),p_{e0}(\theta),1-p_{e0}(\theta)$. This way, the Fisher is computed as $\mbox{FI}=\sum_{i}(1/p_{i})(\partial p_{i}/\partial\theta)^{2}$, where $\{p_{i}\}$ are the four probabilities listed above. This method can enhance the FI only by a factor of $\approx 1.2$ compared to using Eq. (4), despite requiring twice the number of measurements. Therefore, for the rest of the analysis, we will compute Fisher information using Eq. (4).
@@
It is apparent from Fig. 6 that the SCS offers the highest FI. Although the squeezed vacuum is supposed to give a higher theoretical QFI, the state is harder to generate as it requires high squeezing. Highly squeezed states exhibit increased spreading in phase space, making them susceptible to the Kerr effect, which, in turn, hinders quantum-enhanced sensitivity. A viable strategy to mitigate the impact of the Kerr effect involves minimizing the coupling between the cavity and transmon. By reducing this coupling, the undesirable consequences of the Kerr effect can be alleviated, contributing to enhanced quantum sensitivity. Therefore, we discern that with our device, the optimal states are SCS.
Figure 7: Effect of imperfections on the maximum Fisher information obtained using SCS and squeezed vacuum state via full Hamiltonian simulation. The maximum Fisher information from simulations as a function of the average photon number $\bar{n}$ for SCS (solid curves), considering individual imperfections independently. The pink solid curve is the theoretical QFI. For comparison, we also show the results obtained using squeezed vacuum states (dashed curves).
@@
Appendix D Imperfections in the system
We use gradient ascent pulse engineering (GRAPE) pulse to realize the two-state transfer of the protocol. The GRAPE pulses ($\zeta(t)$ and $\epsilon(t)$) are optimized in an ideal system without imperfections. In practice, each element of the system is susceptible to environmental noise, making it crucial to account for the coherence and energy decay rate of both the cavity and transmon qubit. Table 1 presents the corresponding parameters. To account for decoherence, the simulations in the main text (all the sequence) are performed with the following master equation: